# Neural Honeytrace: A Robust Plug-and-Play Watermarking Framework against Model Extraction Attacks

### v1

@@ 3.1 Elements
Channel. According to Shannon’s theorem [34], the channel capacity $C$ can be calculated by:
@@ 2. Why existing triggerable watermarks fail to transmit in adaptive attack scenarios?
However, for Top-1 Attack scenarios, the maximum channel capacity satisfies $C_{max}=\texttt{log}N$. In Fig. 4, $C_{max}=\texttt{log}10\approx 2.3$, according to Eq. 4, the error rate $e\rightarrow 1$, which means the channel is not able to transmit watermark information with a high probability. More importantly, even if the value of $N$ becomes large enough for transmitting watermark information, the channel is exclusive, i.e., only one of the watermark or label information can be transmitted through a single query, resulting in failed transmission.
@@ 4.1 Overview
For Neural Honeytrace, defenders initially select watermarks to be transmitted from the watermark pool (e.g., random watermark, semantic watermark, and composite watermark etc.). Then, the target model extracts the watermark features and query features respectively and makes predictions on the query samples. Neural Honeytrace calculates the similarity of these two features and embeds the similarity into the predictions by mixing the reference logits (target logits of watermarking) and the original logits following the similarity value. Subsequently, Neural Honeytrace establishes a probability-guided label-flipping matrix according to the similarity. By encoding similarity values into the prediction distribution of different queries, this matrix enables multi-step watermarking transmission.
@@ 4.2 Training-free Watermark Embedding
Similarity calculation. After selecting the watermark forms, Neural Honeytrace calculates the distances between input queries and registered watermarks. This process is also performed in the last latent space following the information bottleneck theory. Specifically, the similarity of input query $X$ and registered watermarks can be calculated as:
@@
Similarity embedding. Given the similarity, Neural Honeytrace uses logits-mixing to embed the similarity information into the predictions. Denotes the original logits as $l_{ori}$, the mixed logits can be calculated as:
@@
where $l_{ref}$ is the reference logits of the target watermark class, and $\alpha>1$ establishes a exponentially increasing relationship between the similarity and the activation of watermark targets. By mixing the logits of real samples, Neural Honeytrace reduces the anomaly of the modified logits.
Algorithm 1 Neural Honeytrace
2:OUTPUT: Flipped Logits $l_{flip}$*Watermark Generation**
4:for index $i=1$ to $n$ do
5:     ${W}_{i}\leftarrow{X}_{i}^{k}\cdot{M}+{X}_{i}^{j}\cdot(1-{M})$
6:end for**Similarity Calculation**
7:for index $i=1$ to $n$ do
9:end for
10:if $\text{Max}(\text{SoftMax}(\mathcal{F}({X}_{q})))\geq 0.95$ then
11:     $s\leftarrow s^{2}$
12:end if**Similarity Embedding**
13:${X_{t}}\leftarrow$ Sample($\mathbb{D}$)
14:$l_{mix}\leftarrow(1-s^{\alpha})\cdot\mathcal{F}({X}_{q})+s^{\alpha}\cdot\mathcal{F}({X_{t}})$*Label Flipping**
15:$l_{flip}\leftarrow l_{mix}$
16:if $\text{Bernoulli}(s^{\beta})=1$ then
17:     $l_{flip}\leftarrow\mathcal{F}({X_{t}})+\text{Rand}(\epsilon)$
18:end if
19:return $l_{flip}$
@@ 7 Conclusion, Limitations, and Future Work
In this paper, we propose Neural Honeytrace, a plug-and-play watermarking framework. We first model watermark transmission problem from the perspective of information theory. Based on the theoretical analysis, we propose two watermarking strategies: training-free watermark embedding and multi-step watermark transmission, to achieve training-free watermarking and enhance the robustness of watermarking. Experimental results show that Neural Honeytrace is significantly more robust against adaptive attacks. It is also more flexible and can be adjusted or removed in time if needed.

### v2

@@ 3.1 Elements
Channel. According to Shannon’s theorem [34], the channel capacity $C$ can be calculated using mutual information $I$:
@@ 2. Why existing triggerable watermarks fail to transmit in adaptive attack scenarios?
However, for Top-1 Attack scenarios [23, 37], the maximum channel capacity satisfies $C_{max}=\texttt{log}N$. In Fig. 4, $C_{max}=\texttt{log}10\approx 2.3$, according to Eq. 4, the error rate $e\rightarrow 1$, which means the channel is not able to transmit watermark information with a high probability. Moreover, even if the value of $N$ becomes large enough for transmitting watermark information, the channel is exclusive, i.e., only one of the watermark or label information can be transmitted through a single query, resulting in failed transmission.
@@ 4.1 Overview
For Neural Honeytrace, the watermarking process consists of four steps. Step 1: defenders initially select watermarks to be transmitted from the watermark pool (e.g., random watermark, semantic watermark, and composite watermark etc.). Step 2: the target model extracts the watermark features and query features respectively and makes predictions on the query samples. Step 3: Neural Honeytrace calculates the similarity of these two features and embeds the similarity into the predictions by mixing the reference logits (target logits of watermarking) and the original logits following the similarity value. Step 4: Neural Honeytrace establishes a probability-guided label-flipping matrix according to the similarity. By encoding similarity values into the prediction distribution of different queries, this matrix enables multi-step watermarking transmission, which is robust against adaptive attacks.
@@ 4.2 Training-free Watermark Embedding
Similarity calculation. After selecting the watermark forms, Neural Honeytrace calculates the distances between input queries and registered watermarks. This process is also performed in the last latent space following the information bottleneck theory. Specifically, the similarity of input query $X$ and $N$ registered watermarks can be calculated as:
@@
Similarity embedding. Given the similarity, Neural Honeytrace uses logits-mixing to embed the similarity information into the predictions. Denotes the original logits as $l_{ori}$, the reference logits of the target watermark class as $l_{ref}$, then the mixed logits can be calculated as:
@@
where $\alpha>1$ establishes a exponentially increasing relationship between the similarity and the activation of watermark targets. By mixing the logits of real samples, Neural Honeytrace reduces the anomaly of the modified logits.
@@ 4.3 Multi-step Watermark Transmission
Algorithm 1 Neural Honeytrace
2:OUTPUT: Flipped Logits $l_{flip}$
4:for index $i=1$ to $n$ do
5:     ${W}_{i}\leftarrow{X}_{i}^{k}\cdot{M}+{X}_{i}^{j}\cdot(1-{M})$triangleright Watermark Generation
6:end for
7:for index $i=1$ to $n$ do
9:end for
10:if $\text{Max}(\text{SoftMax}(\mathcal{F}({X}_{q})))\geq 0.95$ then
11:     $s\leftarrow s^{2}$triangleright Similarity Calculation
12:end if
13:${X_{t}}\leftarrow$ Sample($\mathbb{D}$)
14:$l_{mix}\leftarrow(1-s^{\alpha})\cdot\mathcal{F}({X}_{q})+s^{\alpha}\cdot\mathcal{F}({X_{t}})$triangleright Similarity Embedding
15:$l_{flip}\leftarrow l_{mix}$
16:if $\text{Bernoulli}(s^{\beta})=1$ then
17:     $l_{flip}\leftarrow\mathcal{F}({X_{t}})+\text{Rand}(\epsilon)$triangleright Label Flipping
18:end if
19:return $l_{flip}$
@@ 7 Conclusion, Limitations, and Future Work
In this paper, we propose Neural Honeytrace, a robust plug-and-play watermarking framework. We first model watermark transmission problem from the perspective of information theory. Based on the theoretical analysis, we propose two watermarking strategies: training-free watermark embedding and multi-step watermark transmission, to achieve training-free watermarking and enhance the robustness of watermarking. Experimental results show that Neural Honeytrace is significantly more robust against adaptive attacks. It is also more flexible and can be adjusted or removed in time if needed.